# RealCustom++: Representing Images as Real-Word for Real-Time Customization

### v1

@@ 2.1 Text-to-Image Customization
Based on the booming research in text-to-image generation[5, 6, 7, 32, 4] and the recent advancements in diffusion models[33, 34, 35, 36], text-to-image customization, which aims to generate text-driven images for subjects specified by users, has garnered rapidly increasing interest. Existing customization methods follow the pseudo-words paradigm, representing given subjects as pseudo-words and then composing them with the given text for customization. Due to the necessary correspondence between the pseudo-words and the given subjects, existing works are confined to either cumbersome test-time optimization-based methods[2, 3, 10, 13, 15, 16, 12, 37] or optimization-free methods[19, 20, 21, 22, 38, 27] that are trained on object datasets with limited categories.
@@
Optimization-based customization stream. In the optimization-based stream, Textual Inversion[2] was the first to propose representing the given subject through new ”words” in the embedding space of a frozen text-to-image model. DreamBooth[3] instead used a rare token as the pseudo-word and further fine-tuned the entire pre-trained diffusion model for better similarity. Custom Diffusion[10] proposed finding a subset of key parameters for customization generation (i.e., the key and value projection layer in the diffusion model) and only optimizing them. P+[15] introduced an extended textual inversion space (i.e., P+), which learns per-layer pseudo-words for faster convergence. Furthermore, NeTI extends the P+ space to P* space by optimizing to learn separate pseudo-words for each denoising timestep and the denoising U-Net layers. Cones[12] was introduced to optimize residual token embedding for each subject and then compose them through a pre-defined layout. Based on the optimized pseudo-words, many subsequent works[39, 18] further introduce finer image patch features by adding an extra image attention module, significantly improving subject similarity but increasing the risk of overfitting and requiring more regularization mechanisms. The main drawback of this stream is that it requires lengthy optimization time (usually ranging from several minutes to hours) and fine-tuned weights storage for each new subject, bringing excessive computation or memory burden that severely limits their online applications.
@@
Optimization-free customization stream. To address the cumbersome requirement of optimizing for each given subject, the optimization-free customization stream has gained increasing research interest recently. Instead of directly optimizing the pseudo-words to represent the given subjects, this stream typically trains an image encoder to project the given subjects into pseudo-words on object-level datasets (e.g., OpenImages[30], FFHQ[40]). ELITE[19] first proposed a learning-based encoder for fast subject customization, consisting of a global mapping network to encode given subjects into pseudo-words and a local mapping network to maintain subject details. BLIP-Diffusion[20] introduces a multimodal encoder for better subject representation, which includes a multimodal encoder pre-training process and a subject representation learning process. InstantBooth[21] proposed rich visual feature representation by introducing a few adapter layers into pre-trained diffusion models, but their method is limited to a few categories (i.e., human, cat, and dog). Subject Diffusion[41] constructs a large-scale dataset consisting of subject detection bounding boxes, segmentation masks, and text descriptions to train the customization model. However, the generated subject usually remains in the same pose as in the reference images and lacks pose diversity, possibly due to their naive reconstruction training objective. More recently, many works[42, 43, 24, 27] focus on human ID-only customized generation. For example, PhotoMaker[42] proposed stacking ID embeddings as the learned pseudo-words for preserving ID information. InstantID[24] designs an IdentityNet to impose semantic and weak spatial conditions during the training process, integrating facial and landmark images with textual prompts to steer image generation. However, the generated human images also lack subject pose diversity (i.e., the human face can only rigidly face the camera). In summary, these encoder-based works usually show less similarity than optimization-based works and generalize poorly to unseen categories in training.
@@
To conclude, the conflicting semantics and entangled influence scope between foreign pseudo-words and native text limit current methods from achieving both optimal similarity and controllability, hindering real-time open-domain customization.
@@ 2.2 Multiple subject customization
Though most existing text-to-image customization methods focus on single-subject customization, interest in multiple-subject customization has also increased. The task of multiple-subject customization can be divided into two categories: decoupling multiple given subjects in a single reference image (denoted as One2Many) and learning each given subject in its own reference image and then composing them into one customized result (denoted as Many2Many). Existing methods for multiple-subject customization follow the foreign pseudo-words paradigm used in single-subject customization, i.e., using different pseudo-words to represent each subject and then designing elaborate algorithms to achieve decoupling between different subjects.
@@
One2Many: decoupling multiple subjects from a single reference image. Break-a-scene[44] was the first to propose extracting a distinct pseudo-word for each subject by augmenting the input image with masks that indicate the presence of target subjects. The mask is provided by users or generated by a pre-trained segmentation model. Other methods[45, 39] proposed automatic mask generation by using the learned pseudo-words’ corresponding cross-attention mask, either through a fixed threshold or Otsu[46] thresholding. DisenDiff[47] proposes an attention calibration method to distribute attention among different concepts and achieve disentanglement without requiring masks. However, the generated customized results often struggle with the background in the reference images.
@@
Many2Many: composing multiple subjects from multiple reference images. This stream customization methods usually focuses on tackling the inter-confusion between different learned subject pseudo-words. Custom Diffusion[10] proposed joint training on multiple concepts and a constrained optimization to merge concepts. However, the optimization process usually requires over ten reference images for each subject to converge. Mix-of-Show[48] adopts an embedding-decomposed LoRA[49] for each subject’s optimization. A similar idea is adopted in MultiBooth[50], where each subject is bounded by pre-defined bounding boxes to define its specific generation area.
@@
To conclude, existing multiple-subject customization research still lags behind single one, as current methods are typically based on optimized single-subject customization formulas. These methods require over ten reference images for each subject and employ various techniques to address inter-confusion among foreign pseudo-words, making them difficult to apply in real-world scenarios.
@@
Figure 3: Illustration of RealCustom++, which employs a novel “train-inference” decoupled framework: (a) During training, RealCustom++ learns the alignment between vision conditions and all real words in the text. This is achieved by the Cross-layer Cross-Scale Projector (CCP) to robustly and finely extract subject features, and a Curriculum Training Recipe (CTR) that adapts the generated subject to diverse poses and sizes. The projected visual conditions are injected into the diffusion models by extending their textual cross-attention with an additional visual cross-attention in each block. (b) During inference, a novel Adaptive Mask Guidance (AMG) strategy is proposed to customize the generation of the specific target real word (”toy” in this case). This is achieved by incorporating two branches for each generation step: a Guidance Branch that constructs the image guidance mask and a Generation Branch that uses the image guidance mask to keep other subject-irrelevant regions uncontaminated.
@@ 2.4 Curriculum Learning
Curriculum learning[65, 66, 67, 68] is a training strategy that organizes the total training data in an ”easy-to-hard” manner to train a deep learning model, aiming to imitate the way humans learn curricula. The main purpose of curriculum learning is to improve the final model’s performance on target tasks and speed up the training process, which has been widely applied in neural machine translation[69], entity structure relation extraction[70], etc. Existing curriculum learning methods focus on measuring data difficulty[71, 72, 73, 74] or designing specific training schedulers[75, 76, 77, 78]. However, the application of curriculum learning in image generative tasks remains almost unexplored. Our study fills this gap and shows how to organize and prepare the training data to gradually transition the training of a customization task from a naive ”easy” reconstruction objective to a ”hard” re-contextualization objective. Such a training process bridges the discrepancy between training and inference, leading to coherent subject generation in diverse poses and sizes.
@@ 3 Method: RealCustom++
We first briefly introduce the preliminaries in Sec. 3.1. The training and inference paradigms of RealCustom++ are elaborated in detail in Sec. 3.2 and Sec. 3.3, respectively. Finally, we illustrate extending our new paradigm to multiple-subject customization tasks in Sec. 3.4.
@@ 3.1 Preliminaries
3.2 Training Paradigm
Given a reference image $\boldsymbol{x}\in\mathbb{R}^{3\times H\times W}$, where $H$ and $W$ denote the image height and width, respectively. The reference image is then projected into vision condition $\boldsymbol{f_{ci}}\in\mathbb{R}^{n_{\text{image}}\times c_{\text{image}}}$ through the proposed cross-layer cross-scale projector (CCP), where $n_{\text{image}}$ and $c_{\text{image}}$ are the vision condition’s token number and dimension. Afterwards, we extend the textual cross-attention in pre-trained diffusion models with an extra visual cross-attention to incorporate the visual condition $\boldsymbol{f_{ci}}$. Specifically, Eq. 3 is rewritten as:
@@ 3.2.1 Cross-layer Cross-scale Projector
In our conference version[31], subject images are encoded from the last hidden states (deep features) of pretrained image ViT encoders, typically CLIP[29] or DINO[80], a standard implementation widely employed by existing customization methods[19, 21, 43, 24]. Given that these image encoders are pretrained with discriminative tasks (e.g., contrastive loss between paired images and text[29] or self-supervised training[80]), their last hidden states exhibit high semantic richness, providing an advantageous starting point for aligning visual and textual conditions. However, upon further exploration of the model proposed in our conference version across more open-domain unseen subject cases during inference, we identified two key limitations of this prevalent practice:
@@
Figure 5: Illustration of a typical failure when naively concatenating shallow features with deep ones, resulting in a ”copy-paste” problem that disrupts the alignment between text and deep image features.
@@
Cross-scale attention to enrich subject details. As depicted in Figure 4, we introduce an additional high-resolution image stream to extract subject representations with enhanced detail. Specifically, we first resize the subject image to twice the target resolution of the pretrained image encoder (e.g., $768^{2}$ for a ViT image encoder[81] pretrained on $384^{2}$). Subsequently, the high-resolution $768^{2}$ subject images are partitioned into four $384^{2}$ local images, which are then fed into the pretrained image encoder. These local image features are concatenated along their token dimension, yielding a quadruple-length feature $\boldsymbol{f_{\text{high}}}\in\mathbb{R}^{(4\times n_{\text{image}})\times c_{0}}$. Directly using these quadruple-length features significantly increases the training cost (i.e., experimentally, the maximum batch size is limited to 1 when directly feeding these features into the visual cross-attention). Another cross-scale attention is implemented to integrate the high-resolution local features with the original low-resolution global features without increasing the token length:
@@ 3.2.2 Curriculum Training Recipe
where $S_{\text{cur}}$ denotes the current training steps and $S_{\text{total}}$ denotes the total training steps. The terms $r_{\text{min}},r_{\text{max}},r_{\text{cur}}$ denote the minimum, maximum, and current image resize ratios, respectively. A random ratio $r_{\text{sample}}$ is sampled for each training step from a uniform distribution $U(r_{\text{min}},r_{\text{cur}})$. Subsequently, we first resize the reference image to $(384\times r_{\text{sample}})^{2}$ and then randomly crop a $384^{2}$ patch from it. This patch serves as the final reference image, which is fed into the image projector to extract image features. Empirically, we set $r_{\text{min}}=1.0$ and $r_{\text{max}}=\sqrt{10}$, resulting in the most extreme case where the randomly cropped patch constitutes a minimum of 10% of the original reference image’s area.
3.3 Inference Paradigm
As depicted in Figure 3(b), the inference paradigm of RealCustom++ comprises two branches at each generation step: a Guidance Branch (GuiB), where the visual condition is set to None, and a Generation Branch (GenB), where the visual condition corresponds to the given subjects. These two branches are interconnected by our proposed Adaptive Mask Guidance (AMG) strategy. Specifically, given the output $\boldsymbol{z_{t}}$ from the previous step, GuiB first performs a pure text-conditional denoising process to obtain the guidance mask, which is subsequently utilized in GenB.
@@ 3.3.1 Guidance Branch
More spatially accurate image guidance mask: On one side, the textual cross-attention in pre-trained diffusion models primarily captures high-level textual semantic correspondences. Consequently, low-resolution cross-attention maps tend to be more focused and less noisy, while high-resolution cross-attention maps are the opposite. On the other side, self-attention primarily captures low-level per-pixel correspondences, resulting in a high-resolution self-attention map that is more fine-grained and thus more accurate. Therefore, considering their strengths and weaknesses, we propose constructing a more spatially accurate image guidance mask than our conference version[31] by multiplying the low-resolution cross-attention maps of the target native real words (e.g., the real word ”toy” in Figure 3(b)) with the high-resolution self-attention maps in the GuiB. Specifically, we first extract and resize all low-resolution cross-attention maps of the target native real words to the largest map size (e.g., $64\times 64$ in Stable Diffusion XL), denoted as $\boldsymbol{M}_{\text{cross}}\in\mathbb{R}^{64\times 64}$. $\boldsymbol{M}_{\text{cross}}$ is further expanded to 1D as $\boldsymbol{M}_{\text{cross}}\in\mathbb{R}^{(64\times 64)\times 1}$. Then, we extract and resize all high-resolution self-attention maps to $\boldsymbol{M}_{\text{self}}\in\mathbb{R}^{4096\times 4096}$. The multiplied attention map is denoted as $\boldsymbol{M}$:
@@
More temporally stable image guidance mask: Through experimentation, we discovered that as the generation process advances, the image guidance mask tends to converge in the middle diffusion steps and become more scattered later, as illustrated in Figure 8. Consequently, we propose an ”early stop” regularization to stabilize the guidance mask in these later stages. Specifically, given a timestep threshold $T_{\text{stop}}$, we reuse the image guidance mask $\boldsymbol{\hat{M}}_{T_{\text{stop}}}$ for diffusion steps later than $T_{\text{stop}}$. This regularization strategy not only stabilizes the image guidance mask for the latter diffusion steps but also accelerates the customization process, as only a single GenB branch is required after $T_{\text{stop}}$.

### v2

@@ 2.1 Text-to-Image Customization
Based on the booming research in text-to-image generation[5, 6, 7, 32, 4] and the recent advancements in diffusion models[33, 34, 35, 36], text-to-image customization, which aims to generate text-driven images for subjects specified by users, has garnered rapidly increasing interest. Existing customization methods follow the pseudo words paradigm, representing given subjects as pseudo words and then composing them with the given text for customization. Due to the necessary correspondence between the pseudo words and the given subjects, existing works are confined to either cumbersome test-time optimization-based methods[2, 3, 10, 13, 15, 16, 12, 37] or optimization-free methods[19, 20, 21, 22, 38, 27] that are trained on object datasets with limited categories.
@@
Optimization-based customization stream. In the optimization-based stream, Textual Inversion[2] was the first to propose representing the given subject through new ”words” in the embedding space of a frozen text-to-image model. DreamBooth[3] instead used a rare token as the pseudo word and further fine-tuned the entire pre-trained diffusion model for better similarity. Custom Diffusion[10] proposed finding a subset of key parameters for customization generation (i.e., the key and value projection layer in the diffusion model) and only optimizing them. P+[15] introduced an extended textual inversion space (i.e., P+), which learns per-layer pseudo words for faster convergence. Furthermore, NeTI extends the P+ space to P* space by optimizing to learn separate pseudo words for each denoising timestep and the denoising U-Net layers. Cones[12] was introduced to optimize residual token embedding for each subject and then compose them through a pre-defined layout. Based on the optimized pseudo words, many subsequent works[39, 18] further introduce finer image patch features by adding an extra image attention module, significantly improving subject similarity but increasing the risk of overfitting and requiring more regularization mechanisms. The main drawback of this stream is that it requires lengthy optimization time (usually ranging from several minutes to hours) and fine-tuned weights storage for each new subject, bringing excessive computation or memory burden that severely limits their practical applications.
@@
Optimization-free customization stream. To address the cumbersome requirement of optimizing for each given subject, the optimization-free customization stream has gained increasing research interest recently. Instead of directly optimizing the pseudo words to represent the given subjects, this stream typically trains an image encoder to project the given subjects into pseudo words on object-level datasets (e.g., OpenImages[30], FFHQ[40]). ELITE[19] first proposed a learning-based encoder for fast subject customization, consisting of a global mapping network to encode given subjects into pseudo words and a local mapping network to maintain subject details. BLIP-Diffusion[20] introduces a multimodal encoder for better subject representation, which includes a multimodal encoder pre-training process and a subject representation learning process. InstantBooth[21] proposed rich visual feature representation by introducing a few adapter layers into pre-trained diffusion models, but their method is limited to a few categories (i.e., human, cat, and dog). Subject Diffusion[41] constructs a large-scale dataset consisting of subject detection bounding boxes, segmentation masks, and text descriptions to train the customization model. However, the generated subject usually remains in the same pose as in the reference images and lacks pose diversity, possibly due to their naive reconstruction training objective. More recently, many works[42, 43, 24, 27] focus on human ID-only customized generation. For example, PhotoMaker[42] proposed stacking ID embeddings as the learned pseudo words for preserving ID information. InstantID[24] designs an IdentityNet to impose semantic and weak spatial conditions during the training process, integrating facial and landmark images with textual prompts to steer image generation. However, the generated human images also lack subject pose diversity (i.e., the human face can only rigidly face the camera). In summary, these encoder-based works usually show less similarity than optimization-based works and generalize poorly to unseen categories in training.
@@
To conclude, the conflicting semantics and entangled influence scope between foreign pseudo words and native text limit current methods from achieving both optimal similarity and controllability, hindering real-time open-domain customization.
@@ 2.2 Multiple subject customization
Though most existing text-to-image customization methods focus on single-subject customization, interest in multiple-subject customization has also increased. The task of multiple-subject customization can be divided into two categories: decoupling multiple given subjects in a single reference image (denoted as One2Many) and learning each given subject in its own reference image and then composing them into one customized result (denoted as Many2Many). Existing methods for multiple-subject customization follow the foreign pseudo words paradigm used in single-subject customization, i.e., using different pseudo words to represent each subject and then designing elaborate algorithms to achieve decoupling between different subjects.
@@
One2Many: decoupling multiple subjects from a single reference image. Break-a-scene[44] was the first to propose extracting a distinct pseudo word for each subject by augmenting the input image with masks that indicate the presence of target subjects. The mask is provided by users or generated by a pre-trained segmentation model. Other methods[45, 39] proposed automatic mask generation by using the learned pseudo words’ corresponding cross-attention mask, either through a fixed threshold or Otsu[46] thresholding. DisenDiff[47] proposes an attention calibration method to distribute attention among different concepts and achieve disentanglement without requiring masks. However, the generated customized results often struggle with the background in the reference images.
@@
Many2Many: composing multiple subjects from multiple reference images. This stream customization methods usually focuses on tackling the inter-confusion between different learned subject pseudo words. Custom Diffusion[10] proposed joint training on multiple concepts and a constrained optimization to merge concepts. However, the optimization process usually requires over ten reference images for each subject to converge. Mix-of-Show[48] adopts an embedding-decomposed LoRA[49] for each subject’s optimization. A similar idea is adopted in MultiBooth[50], where each subject is bounded by pre-defined bounding boxes to define its specific generation area.
@@
To conclude, existing multiple-subject customization research still lags behind single one, as current methods are typically based on optimized single-subject customization formulas. These methods require over ten reference images for each subject and employ various techniques to address inter-confusion among foreign pseudo words, making them difficult to apply in real-world scenarios.
@@
Figure 3: (a) Illustration of the training framework of RealCustom++, which learns the alignment between vision conditions and all words in the text. This is achieved by the Cross-layer Cross-scale Projector (CCP) to robustly and finely extract subject features, and a Curriculum Training Recipe (CTR) that adapts the generated subject to diverse poses and sizes. The projected visual conditions are injected into the diffusion models by extending their textual cross-attention with another visual cross-attention in each block. (b) The inference framework of RealCustom++ (illustrated at the generation step $t$), which introduces a novel Adaptive Mask Guidance (AMG) to customize the generation of the specific target real word (”toy” in this case). This is achieved by incorporating two branches for each generation step: a Guidance Branch that constructs the image guidance mask and a Generation Branch that uses the image guidance mask to keep other subject-irrelevant regions uncontaminated.
@@ 2.4 Curriculum Learning
Curriculum learning[65, 66, 67, 68] is a training strategy that organizes the total training data in an ”easy-to-hard” manner to train a deep learning model, aiming to imitate the way humans learn curricula. The main purpose of curriculum learning is to improve the final model’s performance on target tasks and speed up the training process, which has been widely applied in neural machine translation[69], entity structure relation extraction[70], etc. Existing curriculum learning methods focus on measuring data difficulty[71] or designing specific training schedulers[72, 73, 74] However, the application of curriculum learning in image generative tasks remains almost unexplored. Our study fills this gap and shows how to organize and prepare the training data to gradually transition the training of a customization task from a naive ”easy” reconstruction objective to a ”hard” re-contextualization objective. Such a training process bridges the discrepancy between training and inference, leading to coherent subject generation in diverse poses and sizes.
@@ 3 Method: RealCustom++
We first briefly introduce the preliminaries in Sec. 3.1. The training and inference framework of RealCustom++ are elaborated in detail in Sec. 3.2 and Sec. 3.3, respectively. Finally, we illustrate extending our new paradigm to multiple-subject customization tasks in Sec. 3.4.
@@ 3.1 Preliminaries
3.2 Training Framework
Given a reference image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the image height and width, respectively. As illustrated in Figure.3(a), the reference image is then projected into vision condition $\boldsymbol{f_{ci}}\in\mathbb{R}^{n_{\text{image}}\times c_{\text{image}}}$ through the proposed cross-layer cross-scale projector (CCP), where $n_{\text{image}}$ and $c_{\text{image}}$ are the vision condition’s token number and dimension. Afterwards, we extend the textual cross-attention in pre-trained diffusion models with another visual cross-attention. Specifically, Eq. 3 is rewritten as:
@@ 3.2.1 Cross-layer Cross-scale Projector
In our conference version[31], subject images are encoded from the last hidden states of pretrained image ViT encoders, typically CLIP[29] or DINO[76], a standard practice in existing customization methods[19, 21, 43, 24]. Given that these image encoders are pretrained with discriminative tasks (e.g., contrastive loss between paired images and text[29] or self-supervised training[76]), their last hidden states exhibit high semantic richness, providing an advantageous starting point for aligning visual and textual conditions. However, upon further exploration of the model proposed in our conference version across more open-domain unseen subject cases during inference, we identified two key limitations of this prevalent practice:
@@
Figure 5: Illustration of a typical failure when naively concatenating shallow and deep features, resulting in a ”copy-paste” problem that disrupts text and image feature alignment.
Cross-scale attention to enrich subject details. As depicted in Figure 4, we introduce an additional high-resolution image stream to extract subject representations with enhanced detail. Specifically, we first resize the subject image to twice the target resolution of the pretrained image encoder (e.g., $768\times 768$ for a ViT image encoder[77] pretrained on $384\times 384$). Subsequently, the high-resolution $768\times 768$ subject images are partitioned into four $384\times 384$ local images, which are then fed into the pretrained image encoder. These local image features are concatenated along their token dimension, yielding a quadruple-length feature $\boldsymbol{f_{\text{high}}}\in\mathbb{R}^{(4\times n_{\text{image}})\times c_{0}}$. Directly using these quadruple-length features significantly increases the training cost (i.e., experimentally, the maximum batch size is limited to 1 when directly feeding these features into the visual cross-attention). Therefore, another cross-scale attention is proposed to integrate the high-resolution local features with the original low-resolution global features without increasing the final token length:
@@ 3.2.2 Curriculum Training Recipe
where $S_{\text{cur}}$ denotes the current training steps and $S_{\text{total}}$ denotes the total training steps. The terms $r_{\text{min}},r_{\text{max}},r_{\text{cur}}$ denote the minimum, maximum, and current image resize ratios, respectively. A random ratio $r_{\text{sample}}$ is sampled for each training step from a uniform distribution $U(r_{\text{min}},r_{\text{cur}})$. Subsequently, we first resize the reference image to $(384\times r_{\text{sample}})^{2}$ and then randomly crop a $384\times 384$ patch from it. This patch serves as the final reference image, which is fed into the projector to extract image features. Empirically, we set $r_{\text{min}}=1.0$ and $r_{\text{max}}=\sqrt{10}$, resulting in the most extreme case where the randomly cropped patch constitutes a minimum of 10% of the original reference image’s area.
3.3 Inference Framework
As depicted in Figure 3(b), the inference framework of RealCustom++ comprises two branches at each generation step: a Guidance Branch, where the visual condition is set to None, and a Generation Branch, where the visual condition corresponds to the given subjects. These two branches are interconnected by our proposed Adaptive Mask Guidance (AMG). Specifically, given the output $\boldsymbol{z_{t}}$ from the previous step, Guidance Branch first performs a pure text-conditional denoising process to obtain the guidance mask, which is subsequently utilized in Generation Branch.
@@ 3.3.1 Guidance Branch
More spatially accurate image guidance mask: On one side, the textual cross-attention in pre-trained diffusion models primarily captures high-level textual semantic correspondences. Consequently, low-resolution cross-attention maps tend to be more focused and less noisy, while high-resolution cross-attention maps are the opposite. On the other side, self-attention primarily captures low-level per-pixel correspondences, resulting in a high-resolution self-attention map that is more fine-grained and thus more accurate. Therefore, considering their strengths and weaknesses, we propose constructing a more spatially accurate image guidance mask than our conference version[31] by multiplying the low-resolution cross-attention maps of the target native real words (e.g., the real word ”toy” in Figure 3 with the high-resolution self-attention maps in the Guidance Branch. Specifically, we first extract and resize all low-resolution cross-attention maps of the target native real words to the largest map size (e.g., $64\times 64$ in Stable Diffusion XL), denoted as $\boldsymbol{M}_{\text{cross}}\in\mathbb{R}^{64\times 64}$. $\boldsymbol{M}_{\text{cross}}$ is further expanded to 1D as $\boldsymbol{M}_{\text{cross}}\in\mathbb{R}^{(64\times 64)\times 1}$. Then, we extract and resize all high-resolution self-attention maps to $\boldsymbol{M}_{\text{self}}\in\mathbb{R}^{4096\times 4096}$. The multiplied attention map is denoted as $\boldsymbol{M}$:
@@
Figure 8: Illustration of the motivation for early stop regularization in mask calculation, where the image guidance mask converges in the middle diffusion steps but scatters in the later steps.
More temporally stable image guidance mask: Through experimentation, we discovered that as the generation process advances, the image guidance mask tends to converge in the middle diffusion steps and become more scattered later, as illustrated in Figure 8. Consequently, we propose an ”early stop” regularization to stabilize the guidance mask in these later stages. Specifically, given a timestep threshold $T_{\text{stop}}$, we reuse the image guidance mask $\boldsymbol{\hat{M}}_{T_{\text{stop}}}$ for diffusion steps later than $T_{\text{stop}}$. This regularization strategy not only stabilizes the image guidance mask for the latter diffusion steps but also accelerates the customization process, as only a single Generation Branch is required after $T_{\text{stop}}$.